Calculate two iterations of Newtons Method to approximate a zero of the function using the given initial guess. (Round yourCalculate two iterations of Newton's Method to approximate a zero of the function using the given initial guess. (Round your answers to three decimal places.)

f(x) = x7 − 7,    x1 = 1.2
Answer #1

외(1)트 치면 그 기는 12 d에 케6 For 1st Rtuation ? 웨 = ? flor) 에i) 조 개 길 t.2- +) 나 () (120 2 (-) ㅏ 3.416) 12 - (20.40) (-0-163) 12 (For an Italation: i-2 직 2 th2) f182) ?!= 1.363. (1.363) ?- 7.11-363) z = 1.363 1-739 44.982 Hz = 1.363. (0.0 387) 33 = 1:32please give thumbs up and comment if you have any questions thankyou